A certain planet spins on its axis 2 a day and orbits the sun 13 a year. Determine the average angular velocity (in rad/s) of the earth as it spins on its axis.


1
Expert's answer
2021-12-20T10:29:04-0500
"\omega=2\frac{2\pi}{24}\frac{1}{3600}=0.000145\frac{rad}{s}"